OpenStudy (anonymous):

\[tan^{-1}(tan\frac{4*\pi}{5})=\]

OpenStudy (anonymous):

I get \[\frac\pi{5}\]

OpenStudy (anonymous):

what did I do wrong?

hartnn (hartnn):

its -pi/5 , right ? \(\tan (\theta - \pi)=\tan \theta\) 4pi/5 - pi = -pi/5 also, \(\tan^{-1}\tan \theta = \theta \)

OpenStudy (anonymous):

oh it's period is pi.....I see

hartnn (hartnn):

yes, period of tan theta = pi. thats why its has same values after every 180 degrees...

OpenStudy (anonymous):

makes sense. Thanks!

hartnn (hartnn):

welcome :)